Find each of the following percentile points and draw the appropriate normal



curve. Complete your procedures.



1.The results of the entrance examination for freshmen are normally distributed



with 𝑥 = 87 and 𝑠 = 12.5. What is the percentile rank of a score of 90?

1
Expert's answer
2022-04-12T13:24:18-0400

"P(X>90)=1-P(Z<\frac{90-87}{12.5})=1-P(Z<0.24)=1-0.5948=0.4052"